### v1

# 00footnotetext:  $\ast$ Corresponding author

Abstract
In this paper, under suitable geometric constraints, we have successfully obtained characterizations for the extremum values of the functional of mixed eigenvalues of the Laplacian on triangles (or trapezoids) in the Euclidean plane $\mathbb{R}^{2}$. More precisely, we prove:
  • For the mixed Dirichlet-Neumann eigenvalue problem of the Laplacian on triangles with fixed area and the Dirichlet boundary condition imposed on non-longest two sides (let us call Type-A mixed eigenvalue problem), the isosceles right triangle with the same area minimizes the first mixed eigenvalue of this eigenvalue problem. This fact can be seen as a Faber-Krahn type inequality for the Type-A first mixed eigenvalue of triangles in $\mathbb{R}^{2}$. As an application, this Faber-Krahn type isoperimetric inequality gives a sharp lower bound estimate for the Type-A first mixed eigenvalue (of triangles) in terms of the area of triangles. Moreover, when this lower bound is attained, a rigidity result can be obtained as well.
  • We also give a lower bound estimate (in terms of the length of the longest side) for the Type-A first mixed eigenvalue of triangles using another approach. This lower bound in some cases is more sharper.
  • As a byproduct, for the mixed Dirichlet-Neumann eigenvalue problem of the Laplacian on right triangles with fixed area and the Dirichlet boundary condition imposed on the longest side (let us call Type-B mixed eigenvalue problem), we also give a characterization for the possible extremum value of the first mixed eigenvalue of Type-B.
  • For the mixed Dirichlet-Neumann eigenvalue problem of the Laplacian on trapezoids in $\mathbb{R}^{2}$ (with an average width $m$ and a height $h$) having the Dirichlet boundary condition imposed on the left and the right sides, we show that the rectangle with width $m$ and height $h$ is a minimizer of the $k$-th mixed eigenvalue functional, provided $\frac{h}{m}\leq\frac{1}{k}$.
  • For the mixed Dirichlet-Neumann eigenvalue problem of the Laplacian on right trapezoids in $\mathbb{R}^{2}$, with fixed area and having the Dirichlet boundary condition imposed on two parallel sides and the perpendicular side, the rectangle with the same area and a length-to-width ratio $2:1$ minimizes the first mixed eigenvalue of this eigenvalue problem.
Besides, we also recalled and proposed several open questions for readers who are interested in the topic of this paper.
1Faculty of Mathematics and Statistics,
Key Laboratory of Applied Mathematics of Hubei Province,
Hubei University, Wuhan 430062, China
2Key Laboratory of Intelligent Sensing System and Security (Hubei University), Ministry of Education
Emails: gchenruifeng@163.com (R. F. Chen), jiner120@163.com (J. Mao)
1 Introduction
The problem of minimizing the first Dirichlet eigenvalue of the Laplacian on domains with fixed measure in the Euclidean $n$-space $\mathbb{R}^{n}$ has a long history, which actually appeared in Lord Rayleigh’s book “The theory of sound” (for example in the edition of 1894) more than 100 years ago. It was in the 1920’s that G. Faber [5] and E. Krahn [9] independently and simultaneously obtained the following classical spectral isoperimetric inequality (i.e., the Faber-Krahn inequality):
Theorem 1.1.
([5, 9]) Assume that $\Omega$ is an arbitrary open set of $\mathbb{R}^{n}$, and let $c$ be a positive number, $\mathbb{B}$ be the Euclidean $n$-ball of volume $c$. Then111 In this paper, by the abuse of notations, $|\cdot|$ also stands for the Hausdorff measure of a given geometric object. So, here $|\Omega|=c$ means that the volume of the open domain $\Omega$ is $c$.
[…]
where $\lambda_{1}(\cdot)$ stands for the first Dirichlet eigenvalue of the Laplacian on a prescribed geometric object.
Speaking in other words, from Theorem 1.1, one easily knows:
  • Among all open domains with fixed volume in $\mathbb{R}^{n}$, the ball of the same volume minimizes the first Dirichlet eigenvalue of the Laplacian.
We refer to [7] for a survey on classical results and open questions about minimization problems concerning the lower eigenvalues of the Laplace operator. By the way, very recently, under the constraint of fixed weighted volume and the radial, concave assumptions for the weighted function, we have successfully improved this classical Faber-Krahn inequality for the Laplacian to the case of Witten-Laplacian by mainly using the rearrangement technique and suitably constructing trial functions (see [3]).
One can ask a similar problem of minimizing the first Dirichlet eigenvalue of the Laplacian on polygons with $n$ sides and fixed area in $\mathbb{R}^{2}$. G. Pólya has given a nice partial answer to this question. In fact, he has proven:
Theorem 1.2.
(G. Pólya, see e.g. [7, 8]) The equilateral triangle has the least eigenvalue among all triangles of given area. The square has the least first eigenvalue among all quadrilaterals of given area.
The main tool in Pólya’s proof of Theorem 1.2 is the so-called Steiner symmetrization (see e.g. [13] for this notion). Curiously, comparing with the case of triangles, the proof for the case of quadrilaterals is a little bit simpler. Indeed, a sequence of three Steiner symmetrizations allows us to transform any quadrilateral into a rectangle. Therefore, it is sufficient to look at the minimization problem for the first Dirichlet eigenvalue of the Laplacian among rectangles. Since the first Dirichlet eigenvalue $\lambda_{1}(\Omega)$ on any rectangular domain $\Omega$ can be calculated directly, it is easy by using the Cauchy-Schwarz inequality to see that a square with the same area $|\Omega|$ can minimize the functional $\Omega\mapsto\lambda_{1}(\Omega)$. However, Pólya’s this idea cannot work any more for polygons with $n$ sides ($n\geq 5$) since in general the Steiner symmetrization increases the number of sides of $n$-gones (pentagons and others). So, a natural and hard question is:
  • Open Problem 1. (see [7, Subsection 3.2]) Prove that the regular $n$-gone has the least first Dirichlet eigenvalue of the Laplacian among all the $n$-gones of given area for $n\geq 5$.
As mentioned by A. Henrot at the end of [7, Subsection 3.2], the classical isoperimetric inequality linking area and length for regular $n$-gones (see e.g. [11, Theorem 5.1]) supports the above conjecture and lets us have the reason that Open Problem 1 might be true. By the way, the Open Problem 1 is known as Pólya-Szegö conjecture, which was formally raised by G. Pólya and G. Szegö in [13, page 158]. Actually, except the isoperimetric inequality linking area and length for regular $n$-gones, many numerical experiments have been performed for small values of $n$, also suggesting the validity of the Pólya-Szegö conjecture.
Laplacian eigenvalues are usually interpreted as frequencies of vibrating membranes. Actually, in the context of eigenvalue problems of the Laplacian, Dirichlet boundary condition (BC) corresponds to the vibration of a membrane with boundary fixed, while Neumann BC indicates a free membrane. Based on this reason, intuitively, (as said in [14]) a mixed Dirichlet-Neumann BC means that the membrane is partially attached. Besides, because of the relation between Dirichlet eigenvalues (of the Laplacian) and Neumann eigenvalues, generally it follows that the larger the attached portion, the higher the frequencies.
The above facts invoke us to consider mixed eigenvalue problems of the Laplacian on planar $n$-gones with $n\geq 3$, i.e. the eigenequaiton of the Laplacian holds in the inner part of $n$-gones, while a Dirichlet-Neumann BC was imposed on the $n$-sides. The mixed Dirichlet-Neumann conditions (in literatures) are also known as the Zaremba problem (see [15]), and can be highly singular, in particular when a transition between the Dirichlet and Neumann conditions happens in the middle of a side of a polygon (see e.g. [2, 10] for details). Based on this reason, in this paper, we investigate a mixed eigenvalue problem of the Laplacian, with imposing Dirichlet conditions on various combinations of the sides of a triangle or a trapezoid, and only treat the cases with transitions at vertices, which pose no problems.
Let $T\subset\mathbb{R}^{2}$ be a planar triangle in the Euclidean plane $\mathbb{R}^{2}$, and let $L$, $M$, $S$ be the lengths of sides of the triangle $T$ such that $L\geq M\geq S$. Denote by $\lambda_{1}^{set}$ the smallest eigenvalue corresponding to the Dirichlet conditions applied to a chosen set of sides – in this setting, e.g., $\lambda_{1}^{LS}$ would correspond to the Dirichlet conditions imposed on the longest and shortest sides. Consider the mixed Dirichlet-Neumann eigenvalue problem of the Laplacian on the triangle $T$ as follows
$\left\{\begin{array}[]{lll}\Delta u+\lambda^{\mathcal{D}}u=0&\mathrm{in}~~T,\\ u=0&\mathrm{on}~~\mathcal{D}\subset\{L,M,S\},\\ \partial_{\nu}u=0&\mathrm{on}~~\partial T\setminus\mathcal{D},\end{array}\right$. (1.1)
where $\mathcal{D}$ can be any combination of the triangle’s sides, and $\nu$ denotes the unit outward normal vector along the set $\partial T\setminus\mathcal{D}$. As usual, $\Delta$ denotes the Laplace operator. For simplicity and convenience, denote by $\lambda=\lambda^{LMS}$ the (pure) Dirichlet eigenvalues and $\mu=\lambda^{\emptyset}$ the (pure) Neumann eigenvalues. It is well-known that the mixed eigenvalue problem (1.1) only has a discrete spectrum, since $T$ is bounded. Moreover, from [14, Section 2], it is not hard to see that eigenvalues in this discrete spectrum can be listed non-decreasingly as follows
$0<\lambda^{\mathcal{D}}_{1}(T)<\lambda^{\mathcal{D}}_{2}(T)\leq\lambda^{\mathcal{D}}_{3}(T)\leq\cdots\uparrow\infty,$
as long as $\mathcal{D}$ is nonempty. When $\mathcal{D}$ is empty (i.e. the purely Neumann case), one has222 Without specification and for convenience, in the sequel we wish to write the $i$-th mixed eigenvalue $\lambda^{\mathcal{D}}_{i}(T)$ as $\lambda^{\mathcal{D}}_{i}$, $i\in\mathbb{Z}_{+}$ with $\mathbb{Z}_{+}$ the set of all positive integers. This convention would be used similarly for other kinds of eigenvalues investigated in this paper. For instance, the $i$-th Neumann eigenvalue $\mu_{i}(T)$ would be simply written as $\mu_{i}$.
$0=\mu_{1}(T)<\mu_{2}(T)\leq\mu_{3}(T)\leq\mu_{4}(T)\leq\cdots\uparrow\infty.$
Besides, all the eigenvalues of (1.1) can be obtained by minimizing the following Rayleigh quotient
$\mathcal{R}[u]=\frac{\int_{T}|\nabla u|^{2}}{\int_{T}u^{2}},$
where $\nabla$ is the gradient operator in $\mathbb{R}^{2}$. Here, for convenience, we have dropped the integral measure in all the integrations of $\mathcal{R}[u]$, and, without specification, this convention would be also used in the sequel. Specially, the first mixed eigenvalue $\lambda^{\mathcal{D}}_{1}(T)$ and the first nonzero Neumann eigenvalue $\mu_{2}(T)$ can be characterized as follows:
$\lambda^{\mathcal{D}}_{1}(T)=\inf\left\{\mathcal{R}[u]\bigg{|}u\in W^{1,2}(T),u=0~\mathrm{on}~\mathcal{D}\right\}$
and
$\mu_{2}(T)=\inf\left\{\mathcal{R}[u]\bigg{|}u\in W^{1,2}(T),\int_{T}u=0\right\},$
where as usual $W^{1,2}(T)$ stands for a Sobolev space (i.e. the completion of the set of smooth functions $C^{\infty}(T)$ under the Sobolev norm $\|u\|_{1,2}:=\int_{T}u^{2}+\int_{T}|\nabla u|^{2}<\infty$). We refer readers to e.g. [1] for an overview of the variational approach.
Nearly ten years ago, B. Siudeja [14] considered the mixed eigenvalue problem (1.1) and gave an order for the mixed Dirichlet-Neumann eigenvalues. In fact, he proved:
Theorem 1.3.
([14]) For any right triangle with the smallest angle satisfying $\pi/6<\alpha<\pi/4$, it holds
$0=\mu_{1}<\lambda_{1}^{S}<\lambda_{1}^{M}<\mu_{2}<\lambda_{1}^{L}<\lambda_{1}^{MS}<\lambda_{1}^{LS}<\lambda_{1}^{LM}<\lambda_{1}.$
When $\alpha=\pi/6$ (half-of-equilateral triangle) we have $\lambda_{1}^{M}=\mu_{2}$, and for $\alpha=\pi/4$ (right isosceles triangle) we have $S=M$ and $\lambda_{1}^{L}=\mu_{2}$. All other inequalities stay sharp in these cases.
Furthermore for arbitrary triangle, it holds
$\min\{\lambda_{1}^{S},\lambda_{1}^{M},\lambda_{1}^{L}\}<\mu_{2}<\lambda_{1}^{MS}<\lambda_{1}^{LS}<\lambda_{1}^{LM}$
as long as the appropriate sides have different lengths. However, it is possible that $\mu_{2}>\lambda_{1}^{L}$ (for any small perturbation of the equilateral triangle) or $\mu_{2}<\lambda_{1}^{M}$ (for right triangles with $\alpha<\pi/6$).
Moreover, he conjectured that all cases missing in Theorem 1.3 are still true:
  • Open Problem 2. (see [14, Section 1]) For an arbitrary triangle, it holds
    $\lambda_{1}^{S}<\lambda_{1}^{M}<\lambda_{1}^{L}<\lambda_{1}^{MS}$
    as long as appropriate sides have different lengths.
From Theorem 1.2, one knows that for the extremum problem
[…]
equilateral triangles of the same area $c$ should be the unique minimizer (up to the congruence). Obviously, in this case, $\mathcal{D}=\{L,M,S\}$. A natural and interesting problem is trying to solve
[…]
if $\mathcal{D}$ is not equal to the set $\{L,M,S\}$ and is not empty.
We have successfully solved the extremum problem (1.2) in the case $\mathcal{D}=\{M,S\}$. More precisely, we have proven:
Theorem 1.4.
Assume that $T$ is an arbitrary triangle in $\mathbb{R}^{2}$, and $T^{\ast}$ is an isosceles right triangle with the same area as $T$, i.e. $|T|=|T^{\ast}|$. Then one has
$\lambda_{1}^{MS}(T)\geq\lambda_{1}^{MS}(T^{\ast}).$ (1.3)
Moreover, the equality in (1.3) can be attained if and only if $T$ is congruent with $T^{\ast}$.
Speaking in other words, from Theorem 1.4, one knows:
  • Among all triangles with fixed area in $\mathbb{R}^{2}$, the isosceles right triangle of the same area minimizes the functional $T\mapsto\lambda_{1}^{MS}(T)$.
As a byproduct, we can also prove:
Theorem 1.5.
Let $T_{R}$ be an arbitrary right triangle in $\mathbb{R}^{2}$, and $T^{\ast}_{R}$ be an isosceles right triangle with the same area as $T_{R}$, i.e. $|T_{R}|=|T^{\ast}_{R}|$. Then $\lambda_{1}^{L}(T_{R})\geq\lambda_{1}^{L}(T^{\ast}_{R})$.
Remark 1.6.
(1) At the place which is right behind the proof of Theorem 1.4 given in Section 2, we have shown that $\lambda_{1}^{MS}(T^{\ast})=\pi^{2}/|T^{\ast}|$. Therefore, the statement of Theorem 1.4 can be rewritten as:
  • For an arbitrary triangle $T$ in $\mathbb{R}^{2}$, one has
    $\lambda_{1}^{MS}(T)\geq\frac{\pi^{2}}{|T|},$ (1.4)
    and the equality can be attained if and only if $T$ is an isosceles right triangle.
(2) Clearly, the assertion of Theorem 1.2 can be seen as a Faber-Krahn type inequality for the first mixed eigenvalue $\lambda^{MS}_{1}(\cdot)$ of triangles in $\mathbb{R}^{2}$. Besides, our conclusion here can also be seen as a complement to the lower bound estimate for the first eigenvalue of the mixed eigenvalue problem (2.1) (considered in [4, Sections 2 and 3]) of the Laplacian on a bounded $C^{2}$-domain in $\mathbb{R}^{n}$, $n\geq 2$.
The meaning of Remark 1.6 is:
  • By applying Theorem 1.2, a lower bound for the first mixed eigenvalue $\lambda^{MS}_{1}(T)$ in terms of the area of a given triangle $T$ can be successfully given. Moreover, an intersting rigidity conclusion can be obtained when the lower bound was achieved.
Is it possible to give interesting lower bounds for $\lambda^{MS}_{1}(T)$ in terms of other important geometric quantities of the triangle $T$? The answer is affirmative. In fact, motivated by P. Freitas’s work [6] of giving upper and lower bounds for the first Dirichlet eigenvalue of a triangle, we can prove the following lower bound estimate.
Theorem 1.7.
For an arbitrary triangle $T$ in $\mathbb{R}^{2}$, we have
$\lambda_{1}^{MS}(T)\geq\frac{4\pi^{2}}{\ell^{2}},$ (1.5)
where $\ell$ is the length of the longest side $L$ of the triangle $T$.
Remark 1.8.
We wish to make a comparison between the estimate (1.4) and the estimate (1.5). Let $\alpha$ be the angle corresponding to the longest side $L$, and let $m$, $s$, $\ell$ be the lengths of sides $S$, $M$ and $L$ of the triangle $T$, respectively. We divide the comparison into two cases:
Case I. When $\alpha\geq\frac{\pi}{2}$, one has
$m^{2}+s^{2}\geq 2ms\geq 2ms\sqrt{2}\sin(\alpha+\frac{\pi}{4}),$
which implies
$m^{2}+s^{2}\geq 2ms(\sin\alpha+\cos\alpha).$
Simplifying the above inequality yields
$m^{2}+s^{2}-2ms\cos\alpha\geq 2ms\sin\alpha,$
which directly implies
$\ell^{2}\geq 4|T|.$
This is to say that when $\alpha\geq\frac{\pi}{2}$ (i.e. $T$ is a right triangle or an obtuse triangle), the lower bound estimate (1.4) is sharper than (1.5). Especially, if $\ell^{2}=4|T|$, then $m=s$ and $\alpha=\pi/4$, which implies that in this setting $T$ should be an isosceles right triangle.
Case II. Since $\alpha$ is the angle corresponding to the longest side $L$, it is not hard to get $\alpha\geq\frac{\pi}{3}$. Now, we discuss the case $\frac{\pi}{3}\leq\alpha<\frac{\pi}{2}$ (i.e. $T$ is an acute triangle).
(1) If furthermore $m=s$ (i.e. $T$ is an isosceles acute triangle), then we first have $\sin(\alpha+\frac{\pi}{4})>\frac{\sqrt{2}}{2}$, which implies
$2s^{2}<2s^{2}\sqrt{2}\sin(\alpha+\frac{\pi}{4}),$
and so
$2s^{2}<2s^{2}(\sin\alpha+\cos\alpha).$
Simplifying the above inequality yields
$2s^{2}-2s^{2}\cos\alpha<4\times\left(\frac{1}{2}s^{2}\sin\alpha\right),$
that is, $\ell^{2}<4|T|$. So, in this setting, the lower bound estimate (1.5) is sharper than (1.4).
(2) If furthermore $m\neq s$, then the relation between $\ell^{2}$ and $4|T|$ is uncertain and deeply depends on the choice of values of $m$, $s$. We wish to give two examples to explain this situation. For instance, considering triangles with the maximal interior angle $\alpha=5\pi/12$, if $m=4$, $s=3$, a simple calculation gives us
$4|T|\approx 23.18,\qquad\ell^{2}\approx 18.7888,$
which means $4|T|>\ell^{2}$ in this setting; if $m=4$, $s=2$, a simple calculation tells us
$4|T|\approx 15.454,\qquad\ell^{2}\approx 17.93,$
which implies $4|T|<\ell^{2}$ in this setting.
In sum, through the above comparison, readers can see that it is meaningful to give estimates (1.4)-(1.5) for the mixed eigenvalue $\lambda^{MS}_{1}(T)$ simultaneously, and these estimates deeply and intuitively reveal the relation between the eigenvalue $\lambda^{MS}_{1}(T)$ and the length of the longest side, the area of the given triangle $T$.
Given a trapezoid in $\mathbb{R}^{2}$, assume that the mean length of the two parallel sides of this trapezoid is $m$ (we call $m$ the mean width or the average width of the trapezoid), and the height is $h$. For Dirichlet eigenvalues of the Laplacian on this given trapezoid, G. Pólya [12] proved:
Theorem 1.9.
([12]) When $\frac{h}{m}\leq\frac{\sqrt{3}}{k}$, $k=1,2,3,\cdots$. A rectangle with an average width $m$ and a height $h$ achieves the minimum value of the $k$-th Dirichlet eigenvalue $\lambda_{k}$.
Inspired by the conclusion of Theorem 1.9, it is natural to consider a similar question for the mixed Dirichlet-Neumann eigenvalue problem of the Laplacian on trapezoids.
Let $\Gamma$ be a trapezoid in $\mathbb{R}^{2}$. Denote the shorter and the longer parallel sides of $\Gamma$ by $P_{1}$, $P_{2}$ respectively. Let the rest sides (i.e. the left and the right legs of $\Gamma$) be denoted by $Q_{1}$, $Q_{2}$. Consider the mixed Dirichlet-Neumann eigenvalue problem of the Laplacian on the trapezoid $\Gamma$ as follows
$\left\{\begin{array}[]{lll}\Delta u+\lambda^{\mathcal{D}}u=0&\mathrm{in}~~\Gamma,\\ u=0&\mathrm{on}~~\mathcal{D}\subset\{P_{1},P_{2},Q_{1},Q_{2}\},\\ \partial_{\nu}u=0&\mathrm{on}~~\partial\Gamma\setminus\mathcal{D},\end{array}\right$. (1.6)
where, similar as before, $\mathcal{D}$ can be any combination of the trapezoid’s sides, and $\nu$ denotes the unit outward normal vector along the set $\partial\Gamma\setminus\mathcal{D}$. Of course, as explained before, in order to avoid singularities arisen, we only treat the cases with transitions at vertices of $\Gamma$. Similar as the case of triangles, it is not hard to see that the mixed eigenvalue problem (1.6) only has a discrete spectrum, and eigenvalues in this discrete spectrum can be listed non-decreasingly as follows
$0<\lambda^{\mathcal{D}}_{1}(\Gamma)<\lambda^{\mathcal{D}}_{2}(\Gamma)\leq\lambda^{\mathcal{D}}_{3}(\Gamma)\leq\cdots\uparrow\infty,$
as long as $\mathcal{D}$ is nonempty. When $\mathcal{D}$ is empty (i.e. the purely Neumann case), one has
$0=\mu_{1}(\Gamma)<\mu_{2}(\Gamma)\leq\mu_{3}(\Gamma)\leq\mu_{4}(\Gamma)\leq\cdots\uparrow\infty.$
By applying the variational principle, one knows that all the eigenvalues of (1.6) can be obtained by minimizing the Rayleigh quotient $\mathcal{R}[u]$ (with $T$ replaced by $\Gamma$), and the first mixed eigenvalue $\lambda^{\mathcal{D}}_{1}(\Gamma)$, the first nonzero Neumann eigenvalue $\mu_{2}(\Gamma)$ can be characterized similarly. Using the convention before, if $\mathcal{D}=\{P_{1},P_{2},Q_{1},Q_{2}\}$, we wish to write $\lambda_{k}^{\mathcal{D}}(\Gamma)=\lambda_{k}^{P_{1}P_{2}Q_{1}Q_{2}}(\Gamma)$ as $\lambda_{k}(\Gamma)$, $k=1,2,\cdots$. Besides, if there is no confusion, $\lambda_{k}(\Gamma)$ can be simplified as $\lambda_{k}$ directly.
For the mixed eigenvalue problem (1.6), we have:
Theorem 1.10.
Assume that $\Gamma\subset\mathbb{R}^{2}$ is a trapezoid with an average width $m$ and a height $h$, and assume that $\Gamma^{\ast}$ is a rectangle with width $m$ and height $h$. If $\frac{h}{m}\leq\frac{1}{k}$, then one has
$\lambda_{k}^{Q_{1}Q_{2}}(\Gamma)\geq\lambda_{k}^{Q_{1}Q_{2}}(\Gamma^{\ast})$
for $k\in\mathbb{Z}_{+}$.
We also have:
Theorem 1.11.
Assume that $\Gamma_{R}$ is a right trapezoid in $\mathbb{R}^{2}$. Denote the upper and lower parallel sides of $\Gamma_{R}$ by $l_{1}$, $l_{2}$ respectively. Let $w_{2}$ be the side of $\Gamma_{R}$, which is perpendicular with parallel sides $l_{1}$, $l_{2}$, and let $w_{1}$ be the last slant side. Assume that $\Gamma_{R}^{\ast}$ is a rectangle with area $|\Gamma_{R}^{\ast}|=|\Gamma_{R}|$ (i.e. $\Gamma_{R}^{\ast}$’s area equals the area $|\Gamma_{R}|$ of $\Gamma_{R}$) and a length-to-width ratio $2:1$. Then we have
$\lambda_{1}^{l_{1}l_{2}w_{1}}(\Gamma_{R})\geq\lambda_{1}^{l_{1}l_{2}w_{1}}(\Gamma_{R}^{\ast}).$
Moreover, the equality can be achieved if and only if $\Gamma_{R}$ is congruent with $\Gamma_{R}^{\ast}$.
The paper is organized as follows. In Section 2, the proofs of the conclusions for triangles will be given. The case of trapezoids will be dealt with in Section 3.
2 The case of triangles
We give the proof of our first main result.
Proof of Theorem 1.4.
For an arbitrary triangle $T\subset\mathbb{R}^{2}$, we use red solid line segments to represent the sides assigned the Dirichlet boundary condition, and the black dashed line segment to represent the side assigned the Neumann boundary condition. Assume that $u$ is an eigenfunction corresponding to $\lambda_{1}^{MS}(T)$. Now, we fold the triangle $T$ along the longest line $L$, and denote the resulting quadrilateral by $Q$. One knows that the area $|Q|$ of $Q$ should be twice of $|T|$, as shown in the Figure 1.
Figure 1: The quadrilateral $Q$ generated by the triangle $T$ through symmetrization
Construct a function $f$ defined on $Q$ as follows
$f(x)=\left\{\begin{array}[]{ll}u(x)&\mathrm{if}~x\in T,\\ u(y)&\mathrm{if}~x\in Q\setminus T,\end{array}\right$.
where $y$ is the symmetric point of $x$ with respect to the side $BC$. Due to the symmetry of $f$, one has
$\int_{T}|\nabla u|^{2}=\int_{Q\setminus T}|\nabla f|^{2}$
and
$\int_{T}|u|^{2}=\int_{Q\setminus T}|f|^{2}.$
Since $f$ can be regarded as a trial function for the first Dirichlet eigenvalue of the Laplacian on $Q$, one can get
$\lambda_{1}^{MS}(T)$ $=$ $\frac{\int_{T}|\nabla u|^{2}}{\int_{T}u^{2}}=\frac{\int_{T}|\nabla u|^{2}+\int_{Q\setminus T}|\nabla f|^{2}}{\int_{T}u^{2}+\int_{Q\setminus T}f^{2}}$ (2.2)
$=$ $\frac{\int_{Q}|\nabla f|^{2}}{\int_{Q}f^{2}}\geq\lambda_{1}(Q).$
By (2.2) and Theorem 1.2 , one has
$\lambda_{1}^{MS}(T)\geq\lambda_{1}(Q)\geq\lambda_{1}(S_{Q})$ (2.3)
where $S_{Q}\subset\mathbb{R}^{2}$ is a square having the same area as $Q$, i.e. $|S_{Q}|=|Q|$. Since an eigenfunction $h$ of the first Dirichlet eigenvalue of the Laplacian on a square is symmetric with respect to the diagonal of this square, one can divide the above square $S_{Q}$ along its diagonal into two parts of equal area (i.e. two congruent isosceles right triangles). Let one of these isosceles right triangles be as $T^{\ast}$. Denote by the function $h$ restricted on $T^{\ast}$ as $g$, and then it is clear that $g$ takes the value 0 on the two perpendicular sides of triangle $T^{\ast}$. By the variational principle of the mixed Dirichlet-Neumann eigenvalue problem (1.1), it is true that the function $g$ can be served as a trial function for the eigenvalue $\lambda_{1}^{MS}(T^{\ast})$. Due to the symmetry of $h$ with respect to the diagonal of $S_{Q}$, it follows that
$\lambda_{1}(S_{Q})$ $=$ $\frac{\int_{S_{Q}}|\nabla h|^{2}}{\int_{S_{Q}}h^{2}}=\frac{\int_{T^{\ast}}|\nabla g|^{2}+\int_{S_{Q}\setminus T^{\ast}}|\nabla h|^{2}}{\int_{T^{\ast}}g^{2}+\int_{S_{Q}\setminus T^{\ast}}h^{2}}$ (2.4)
$=$ $\frac{\int_{T^{\ast}}|\nabla g|^{2}}{\int_{T^{\ast}}g^{2}}\geq\lambda_{1}^{MS}(T^{\ast}).$
Together with (2.3), it yields $\lambda_{1}^{MS}(T)\geq\lambda_{1}^{MS}(T^{\ast})$. Moreover, when $\lambda_{1}^{MS}(T)=\lambda_{1}^{MS}(T^{\ast})$, one directly gets $\lambda_{1}(Q)=\lambda_{1}(S_{Q})$, which by Theorem 1.2 implies that $Q$ is congruent with the square $S_{Q}$. Since $Q$ is generated by folding the triangle $T$ along its side $L$, one knows in this situation that $T$ must be congruent with $T^{\ast}$. This completes the proof of Theorem 1.4. ∎
Using a similar argument to the derivation of (2.2)-(2.3), it is not hard to see that
$\lambda^{MS}_{1}(T^{\ast})\geq\lambda_{1}(S_{Q}).$
Together with (2.4), one has
$\lambda^{MS}_{1}(T^{\ast})=\lambda_{1}(S_{Q})=\frac{2\pi^{2}}{|S_{Q}|}=\frac{\pi^{2}}{|T^{\ast}|}.$
The proof of the second main result can be given as follows:
Proof of Theorem 1.5.
We take the right triangle $T_{R}:=\triangle ABC$ and reflect it along its two perpendicular sides to form a rhombus $\mathfrak{R}$, as shown in Figure 2 below.
Figure 2: The rhombus $\mathfrak{R}$ formed from the right triangle $\triangle ABC$ by reflection
Let $v$ be the characteristic function corresponding to $\lambda_{1}^{L}(T_{R})$. Construct a function $\phi$ as follows:
$\phi(x)=\left\{\begin{array}[]{llll}v(x)&\mathrm{if}~x\in T_{R},\\ v(y)&\mathrm{if}~x\in\triangle BCD,\\ v(z)&\mathrm{if}~x\in\triangle ACE,\\ v(w)&\mathrm{if}~x\in\triangle DCE,\end{array}\right$.
where $y$ is the symmetric point of $x$ with respect to the side $BC$, $z$ is the symmetric point of $x$ with respect to $AD$, and $w$ is the centrally symmetric point of $x$ with respect to the vertex $C$. It is easy to see that $\phi=0$ on the boundary of the rhombus $\mathfrak{R}$, which implies that $\phi$ can be used as a trial function for the first Dirichlet eigenvalue $\lambda_{1}(\mathfrak{R})$. Using a similar argument to that in the proof of Theorem 1.4, particularly the idea of deriving (2.2), one can obtain
$\lambda^{L}_{1}(T_{R})\geq\lambda_{1}(\mathfrak{R}).$
Together with Theorem 1.2, it yields
$\lambda^{L}_{1}(T_{R})\geq\lambda_{1}(\mathfrak{R})\geq\lambda_{1}(S_{\mathfrak{R}})=\frac{2\pi^{2}}{|S_{\mathfrak{R}}|}=\frac{\pi^{2}}{2|T_{R}|},$
where $S_{\mathfrak{R}}$ is a square having the same area as $\mathfrak{R}$, i.e. $|S_{\mathfrak{R}}|=|\mathfrak{R}|$. By [1] or [14, Theorem 1.1] (i.e. Theorem 1.3 here), one knows
$\lambda_{1}^{L}(T_{R}^{\ast})=\mu_{2}(T_{R}^{\ast})=\frac{\pi^{2}}{2|T_{R}^{\ast}|}=\frac{\pi^{2}}{2|T_{R}|},$
which implies the assertion of Theorem 1.5. ∎
Using a similar idea to that shown in [6], we have:
Proof of Theorem 1.7.
As shown in Figure 1, for an arbitrary triangle $T:=\triangle ABC\subset\mathbb{R}^{2}$, we use red solid line segments to represent the sides assigned the Dirichlet boundary condition, and the black dashed line segment to represent the side assigned the Neumann boundary condition. Assume that $D$ is the point on the line segment $BC$ (of $T$) such that $AD$ is perpendicular with $BC$. Let us enclose the entire triangle $T$ by a rectangular box $GBCH$, as shown in Figure 3 below.
Figure 3: The rectangle GBCHGBCH constructed to enclose the triangle $T=\triangle ABC$
Let $u$ be a e characteristic function corresponding to the first mixed eigenvalue $\lambda_{1}^{MS}(T)$. Construct a function $\phi$ defined on the rectangle $GBCH$ as follows:
$\phi=\left\{\begin{array}[]{lll}u&\mathrm{on}~~T,\\ u_{13}&\mathrm{on}~~\triangle GAB,\\ u_{24}&\mathrm{on}~~\triangle AHC,\end{array}\right$.
where $u_{13}$ is the function obtained from $u$ by rotating the triangle $\triangle ABD$ with respect to the midpoint of its largest side until it coincides with $\triangle GAB$, and then changing its sign. Similarly, we can define the function $u_{24}$ from $u$ on $\triangle ACD$.
Let the length of $AD$ be $h$. For the cylinder $\mathcal{C}$ which is generated by gluing together two sides $BG$, CHCH of the rectangle $GBCH$, consider the Neumann eigenvalue problem of the Laplacian on $\mathcal{C}$, and then a direct calculation gives us
$\mu_{2}(\mathcal{C})=\min\left\{\frac{\pi^{2}}{h^{2}},\frac{4\pi^{2}}{\ell^{2}}\right\},$ (2.7)
where $\mu_{2}(\mathcal{C})$ is the nonzero Neumann eigenvalue of $\mathcal{C}$. By using the basic properties of triangles, one has $0<h\leq\sqrt{3}\ell/2$, and moreover $h=\sqrt{3}\ell/2$ if and only if $T$ is an equilateral triangle. This fact, together with (2.7), implies
$\mu_{2}(\mathcal{C})=\frac{4\pi^{2}}{\ell^{2}}.$
Due to the definition of $\phi$, it is not hard to see $\int_{\mathcal{C}}\phi=0$, which implies that $\phi$ can be regarded as a trial function for $\mu_{2}(\mathcal{C})$. Hence, we can obtain
$\lambda_{1}^{MS}(T)$ $=$ $\frac{\int_{T}|\nabla u|^{2}}{\int_{T}u^{2}}=\frac{\int_{\triangle ABD}|\nabla u|^{2}+\int_{\triangle ACD}|\nabla u|^{2}}{\int_{\triangle ABD}u^{2}+\int_{\triangle ACD}u^{2}}$
$=$ $\frac{\int_{\triangle ABD}|\nabla u|^{2}+\int_{\triangle ACD}|\nabla u|^{2}+\int_{\triangle AGB}|\nabla u_{13}|^{2}+\int_{\triangle ACH}|\nabla u_{24}|^{2}}{\int_{\triangle ABD}u^{2}+\int_{\triangle ACD}u^{2}+\int_{\triangle AGB}u_{13}^{2}+\int_{\triangle ACH}u_{24}^{2}}$
$=$ $\frac{\int_{\mathcal{C}}|\nabla\phi|^{2}}{\int_{\mathcal{C}}\phi^{2}}$
$\geq$ $\mu_{2}(\mathcal{C})=\frac{4\pi^{2}}{\ell^{2}},$
which completes the proof of Theorem 1.7. ∎
3 The case of trapezoids
We first give the proof of Theorem 1.10. However, the following two facts are needed.
Lemma 3.1.
Let $\Gamma$ be a trapezoid in $\mathbb{R}^{2}$, and rotate $\Gamma$ around the midpoints of its two parallel sides m1m-1 times. Denote these m1m-1 copies of $\Gamma$, also including itself, by $\Gamma=\Omega_{1}$, $\Omega_{2},\cdots,\Omega_{m}$, and moreover, supplement the set $\Omega_{1}\cup\Omega_{2}\cup\cdots\cup\Omega_{m}$ to form a rectangle $\Omega$, as shown in Figure 4. Denote the upper and lower sides of $\Omega$ by $P^{1}$ and $P^{2}$, and the left and right sides by $Q^{1}$ and $Q^{2}$. Then one has
$\lambda_{k}^{Q_{1}Q_{2}}(\Gamma)\geq\lambda_{km}^{Q^{1}Q^{2}}(\Omega),\qquad k=1,2,3,\cdots,$
where, as assumed in (1.6), $Q_{1}$, $Q_{2}$ are the left and the right sides of $\Gamma$.
Figure 4: A rectangle formed by rotating the trapezoid $\Gamma\subset\mathbb{R}^{2}$ and making a supplement
Proof.
Assume first that all the eigenvalues of $\Omega_{1},\Omega_{2},\cdots,\Omega_{m}$ are arranged in ascending order as $v_{1}\leq v_{2}\leq v_{3}\leq\cdots$. Take $\phi_{1},\phi_{2},\cdots,\phi_{k-1}$ to be the k1k-1 corresponding eigenfunctions which are mutually $L^{2}(\Omega)$ orthogonal. For $j=1,2,\cdots,k$, let $\psi_{j}:\Omega\mapsto\mathbb{R}$ be an eigenfunction of $v_{j}$ when restricted to the appropriate subdomain, and identically zero otherwise. Then $\psi_{j}\in W^{1,2}_{0}(\Omega)$. It is not hard to see that $\psi_{j}$ is mutually orthogonal in $L^{2}(\Omega)$. By the fundamental theory of Linear Algebra, there exists a set of nonzero numbers $\alpha_{1},\alpha_{2},\cdots,\alpha_{k}$ such that
$\sum\limits_{j=1}^{k}\alpha_{j}\int_{\Omega}\psi_{j}\phi_{l}=0,\qquad l=1,2,\cdots,k-1.$
Define a function $f$ as follows
$f:=\sum\limits_{j=1}^{k}\alpha_{j}\psi_{j}.$
It is easy to see that $f$ is $L^{2}(\Omega)$ orthogonal to $\phi_{1},\phi_{2},\cdots,\phi_{k-1}$, which implies that $f$ can be used as a trial function for $\lambda_{k}^{Q^{1}Q^{2}}(\Omega)$. Therefore, we have
$\lambda_{k}^{Q^{1}Q^{2}}(\Omega)\int_{\Omega}f^{2}\leq\int_{\Omega}|\nabla f|^{2}=\sum\limits_{j=1}^{k}\alpha_{j}^{2}v_{j}\leq v_{k}\int_{\Omega}f^{2},$
which directly implies
$\lambda_{k}^{Q^{1}Q^{2}}(\Omega)\leq v_{k}.$
Since each $\Omega_{i}$, $i=1,2,\cdots,m$, is congruent with $\Gamma$, one knows that $\lambda_{k}^{Q_{1}Q_{2}}(\Gamma)$ will be repeated $m$ times, and then we can separately arrange all the mixed eigenvalues of $\Omega_{1}\cup\Omega_{2}\cup\Omega_{3}\cup\cdots\cup\Omega_{m}$, $\Omega$ in the following way (i.e. in the first line, all the mixed eigenvalues of $\Omega_{i}$, $i=1,2,\cdots,m$, have been listed non-decreasingly, while in the second line, all the mixed eigenvalues of the rectangle $\Omega$ of the same type have been listed non-decreasingly)
$\bigg{\{}\begin{array}[]{l}\lambda_{1}^{Q_{1}Q_{2}}(\Gamma),\lambda_{1}^{Q_{1}Q_{2}}(\Gamma),\cdots,\lambda_{1}^{Q_{1}Q_{2}}(\Gamma),\lambda_{2}^{Q_{1}Q_{2}}(\Gamma),\lambda_{2}^{Q_{1}Q_{2}}(\Gamma),\cdots,\lambda_{2}^{Q_{1}Q_{2}}(\Gamma),\lambda_{3}^{Q_{1}Q_{2}}(\Gamma),\cdots\\ \lambda_{1}^{Q^{1}Q^{2}}(\Omega),\lambda_{2}^{Q^{1}Q^{2}}(\Omega),\cdots,\lambda_{m}^{Q^{1}Q^{2}}(\Omega),\lambda_{m+1}^{Q^{1}Q^{2}}(\Omega),\lambda_{m+2}^{Q^{1}Q^{2}}(\Omega),\cdots,\lambda_{2m}^{Q^{1}Q^{2}}(\Omega),\lambda_{2m+1}^{Q^{1}Q^{2}}(\Omega),\cdots\end{array}$
which implies the assertion of Lemma 3.1 by making a direct comparison between corresponding eigenvalues of the above two sequences. ∎
Lemma 3.2.
For a rectangle $D$ with length $a$ and width $b$, we still use $P^{1}$, $P^{2}$ to represent the upper and lower sides, and $Q^{1}$, $Q^{2}$ to represent the left and right sides. Then it follows
$\lambda_{k}^{Q^{1}Q^{2}}(D)\geq\frac{\pi^{2}}{b^{2}}M\left(\frac{b}{a}k\right),$
where $M(x)$ is the inverse function of the function $x=\sum\limits_{j=0}^{\infty}(y-j^{2})_{+}^{\frac{1}{2}}$, $y\geq 0$. Besides, here an abbreviation
$z_{+}:=\frac{z+|z|}{2}=\left\{\begin{array}[]{ll}z,&when~z\geq 0\\ 0,&when~z<0\end{array}\right$.
has been used.
Remark 3.3.
From [12, p. 426], one would see that $M(x)$ is a strictly increasing continuous function.
Proof.
Assume that $N$ is the number of eigenvalues whose values are less or equal to $\lambda_{k}^{Q^{1}Q^{2}}(D)$. It is not hard to see that $k\leq N$. Using a method similar with calculating all the Dirichlet eigenvalues of the Laplacian on rectangles, one can obtain all the mixed eigenvalues $\lambda_{l}^{Q^{1}Q^{2}}(D)$ of the Laplacian on the rectangle $D$ as follows
$\pi^{2}\left(\frac{i^{2}}{a^{2}}+\frac{j^{2}}{b^{2}}\right),\qquad i=1,2,3,\cdots,~~j=0,1,2,\cdots.$
Solving the inequality
$\pi^{2}\left(\frac{i^{2}}{a^{2}}+\frac{j^{2}}{b^{2}}\right)\leq\lambda_{k}^{Q^{1}Q^{2}}(D),$
which implies
$i\leq\frac{a}{b}\left(\frac{b^{2}\lambda_{k}^{Q^{1}Q^{2}}(D)}{\pi^{2}}-j^{2}\right)_{+}^{\frac{1}{2}}.$
Hence, we have
$k\leq N=\sum\limits_{j=0}^{\infty}\left[\frac{a}{b}\left(\frac{b^{2}\lambda_{k}^{Q^{1}Q^{2}}(D)}{\pi^{2}}-j^{2}\right)_{+}^{\frac{1}{2}}\right].$
Then, Using the definition of the function $M(x)$ and its monotonicity, we further get
$\lambda_{k}^{Q^{1}Q^{2}}(D)\geq\frac{\pi^{2}}{b^{2}}M\left(\frac{b}{a}k\right).$
This completes the proof of Lemma 3.2. ∎
Now, we have:
Proof of Theorem 1.10.
Let $\widetilde{\Gamma}^{\ast}$ be a rectangle that contains $n$ copies of $\Gamma$, with its average width chosen to be $mn+c$ and its height still $h$, where $c$ depends only on the selection of $\Gamma$ and is independent of $n$. By Lemma 3.1, one has
[…]
By Lemma 3.2, one gets
$\lambda_{kn}^{Q^{1}Q^{2}}(\widetilde{\Gamma}^{\ast})\geq\frac{\pi^{2}}{h^{2}}M\left(\frac{hkn}{mn+c}\right).$ (3.3)
Combining (3.2) and (3.3), letting $n\rightarrow+\infty$, and using the continuity of $M(x)$, we can obtain
$\lambda_{k}^{Q_{1}Q_{2}}(\Gamma)\geq\frac{\pi^{2}}{h^{2}}M\left(\frac{hk}{m}\right).$
By the definition of $M(x)$ and from [12, pp. 426-427], one knows that when $0\leq x\leq 1$, $M(x)=x^{2}$. Therefore, if $\frac{h}{m}\leq\frac{1}{k}$ (i.e. $0<\frac{h}{mk}\leq 1$), we have
$\lambda_{k}^{Q_{1}Q_{2}}(\Gamma)\geq\frac{\pi^{2}}{h^{2}}\frac{k^{2}h^{2}}{m^{2}}=\frac{\pi^{2}k^{2}}{m^{2}}.$ (3.4)
Now, we divide the proof into two cases as follows:
Case I. When $k=1$, one knows
$\lambda_{1}^{Q^{1}Q^{2}}(\Gamma^{\ast})=\frac{\pi^{2}}{m^{2}},$
with $\Gamma^{\ast}$ the rectangle of width $m$ and height $h$ described in Theorem 1.10, and then from (3.4), it follows
$\lambda_{1}^{Q_{1}Q_{2}}(\Gamma)\geq\lambda_{1}^{Q^{1}Q^{2}}(\Gamma^{\ast}).$
Case II. When $k\geq 2$, it is easy to have
$\frac{h^{2}}{m^{2}}\leq\frac{1}{k^{2}}<\frac{1}{k^{2}-1}.$ (3.5)
As we know, all the mixed eigenvalues $\lambda_{l}^{Q^{1}Q^{2}}(\Gamma^{\ast})$ of the rectangle $\Gamma^{\ast}$ are
$\pi^{2}\left(\frac{i^{2}}{m^{2}}+\frac{h^{2}}{b^{2}}\right),\qquad i=1,2,3,\cdots,~~j=0,1,2,\cdots.$
Together with (3.5), we have
$\pi^{2}\left(\frac{i^{2}}{m^{2}}+\frac{h^{2}}{b^{2}}\right)=\frac{\pi^{2}}{m^{2}}\left(i^{2}+j^{2}\cdot\frac{m^{2}}{h^{2}}\right)>\frac{\pi^{2}}{m^{2}}\left[i^{2}+j^{2}\cdot(k^{2}-1)\right].$
Hence, in our setting here, the $k$-th mixed eigenvalue $\lambda_{k}^{Q^{1}Q^{2}}(\Gamma^{\ast})$ should be attained when $i=k$, $j=0$, that is, $\lambda_{k}^{Q^{1}Q^{2}}(\Gamma^{\ast})=\pi^{2}k^{2}/m^{2}$. Together with (3.4), one has
$\lambda_{k}^{Q_{1}Q_{2}}(\Gamma)\geq\lambda_{k}^{Q^{1}Q^{2}}(\Gamma^{\ast}).$
In sum, under the assumption that $\frac{h}{m}\leq\frac{1}{k}$, we always have $\lambda_{k}^{Q_{1}Q_{2}}(\Gamma)\geq\lambda_{k}^{Q^{1}Q^{2}}(\Gamma^{\ast})$, which completes the proof of Theorem 1.10. ∎
At the end, we wish to mainly use the idea of proving Theorem 1.4 to get our last conclusion.
Proof of Theorem 1.11.
We fold the right trapezoid $\Gamma_{R}$ along the right angle side $w_{2}$ to transform it into an isosceles trapezoid $I_{s}$, such that $|I_{s}|=2|\Gamma_{R}|$. Let $v$ be the eigenfunction corresponding to the mixed eigenvalue $\lambda_{1}^{l_{1}l_{2}w_{1}}(\Gamma_{R})$, and construct a function $\rho$ as follows
$\rho(x)=\left\{\begin{array}[]{ll}v(x),&\mathrm{if}~x\in\Gamma_{R},\\ v(y),&\mathrm{if}~x\in I_{s}\setminus\Gamma_{R},\end{array}\right$.
where $y$ is the symmetric point of $x$ with respect to $w_{2}$. It is easy to check that $\rho|_{\partial I_{s}}=0$, which implies that $\rho$ can be regarded as a trial function for the first Dirichlet eigenvalue $\lambda_{1}(I_{s})$ of the Laplacian on $I_{s}$. Based on the symmetry of $\rho$ and using the idea in the proof of Theorem 1.4, we can obtain
$\lambda_{1}^{l_{1}l_{2}w_{1}}(\Gamma_{R})\geq\lambda_{1}(I_{s}).$
By Theorem 1.2, the above inequality can be improved as follows
$\lambda_{1}^{l_{1}l_{2}w_{1}}(\Gamma_{R})\geq\lambda_{1}(I_{s})\geq\lambda_{1}(S_{I_{s}})=\frac{2\pi^{2}}{|I_{s}|}=\frac{\pi^{2}}{|\Gamma_{R}|},$ (3.7)
where $S_{I_{s}}$ denotes a square with area equal to $|I_{s}|$. Moreover, the equality symbol in the second inequality of (3.7) can be achieved if and only if the isosceles trapezoid $I_{s}$ is congruent with the square $S_{I_{s}}$. Since the right hand side of (3.7) can be rewritten as
$\pi^{2}\left(\frac{1}{2|\Gamma_{R}|}+\frac{1}{2|\Gamma_{R}|}\right)=\pi^{2}\left(\frac{1}{2|\Gamma_{R}|}+\frac{1}{4\left(\frac{\sqrt{2|\Gamma_{R}|}}{2}\right)^{2}}\right),$
which is exactly the first mixed eigenvalue $\lambda_{1}^{l_{1}l_{2}w_{1}}(\Gamma_{R}^{\ast})$ of a rectangle $\Gamma_{R}^{\ast}$ with area $|\Gamma_{R}^{\ast}|=|\Gamma_{R}|$ and a length-to-width ratio $2:1$ (see e.g. [1] for this fact), we further have
$\lambda_{1}^{l_{1}l_{2}w_{1}}(\Gamma_{R})\geq\lambda_{1}^{l_{1}l_{2}w_{1}}(\Gamma_{R}^{\ast}).$ (3.8)
By the way, one can use a line segment (joining two midpoints of a pair of parallel sides of the square $S_{I_{s}}$) to divide $S_{I_{s}}$ into two equal parts, and clearly each part is congruent with the rectangle $\Gamma_{R}^{\ast}$. Based on the symmetric property of an eigenfunction of the first Dirichlet Laplacian eigenvalue $\lambda_{1}(S_{I_{s}})$ with respect to the line segment, and using an argument similar to the one shown the fact $\lambda^{MS}_{1}(T^{\ast})=\lambda_{1}(S_{Q})$ in Section 2, it is not hard to get $\lambda_{1}(S_{I_{s}})=\lambda_{1}^{l_{1}l_{2}w_{1}}(\Gamma_{R}^{\ast})$.
If the equality in (3.8) holds, then from the above argument one knows $\lambda_{1}(I_{s})=\lambda_{1}(S_{I_{s}})$, which by Theorem 1.2 implies that $I_{s}$ is congruent with the square $S_{I_{s}}$. Since $I_{s}$ is generated by folding $\Gamma_{R}$ along its right angle side $w_{2}$, one knows in this situation that $\Gamma_{R}$ must be congruent with $\Gamma_{R}^{\ast}$. This completes the proof of Theorem 1.11. ∎
Acknowledgments
This research was supported in part by the NSF of China (Grant Nos. 11801496 and 11926352), the Fok Ying-Tung Education Foundation (China), Hubei Key Laboratory of Applied Mathematics (Hubei University), and Key Laboratory of Intelligent Sensing System and Security (Hubei University), Ministry of Education.

### v2

@@
 Moreover, when this lower bound is attained, a rigidity result can be obtained as well.
-  • We also give a lower bound estimate (in terms of the length of the longest side) for the Type-A first mixed eigenvalue of triangles using another approach.
-This lower bound in some cases is more sharper.
+  • We also give a lower bound estimate (in terms of the length of the longest side or the length of the height of the longest side) for the Type-A first mixed eigenvalue of triangles using another approach.
   • As a byproduct, for the mixed Dirichlet-Neumann eigenvalue problem of the Laplacian on right triangles with fixed area and the Dirichlet boundary condition imposed on the longest side (let us call Type-B mixed eigenvalue problem), we also give a characterization for the possible extremum value of the first mixed eigenvalue of Type-B.
@@
 Based on this reason, in this paper, we investigate a mixed eigenvalue problem of the Laplacian, with imposing Dirichlet conditions on various combinations of the sides of a triangle or a trapezoid, and only treat the cases with transitions at vertices, which pose no problems.
-Let $T\subset\mathbb{R}^{2}$ be a planar triangle in the Euclidean plane $\mathbb{R}^{2}$, and let $L$, $M$, $S$ be the lengths of sides of the triangle $T$ such that $L\geq M\geq S$.
+Let $T\subset\mathbb{R}^{2}$ be a planar triangle in the Euclidean plane $\mathbb{R}^{2}$, and let $L$, $M$, $S$ be the sides of the triangle $T$ such that their lengths satisfy $\mathcal{L}(L)\geq\mathcal{L}(M)\geq\mathcal{L}(S)$.
 Denote by $\lambda_{1}^{set}$ the smallest eigenvalue corresponding to the Dirichlet conditions applied to a chosen set of sides – in this setting, e.g., $\lambda_{1}^{LS}$ would correspond to the Dirichlet conditions imposed on the longest and shortest sides.
@@
 For an arbitrary triangle $T$ in $\mathbb{R}^{2}$, we have
-$\lambda_{1}^{MS}(T)\geq\frac{4\pi^{2}}{\ell^{2}},$ (1.5)
-where $\ell$ is the length of the longest side $L$ of the triangle $T$.
+$\lambda_{1}^{MS}(T)\geq\min\left\{\frac{\pi^{2}}{h^{2}},\frac{4\pi^{2}}{\ell^{2}}\right\},$ (1.5)
+where $\ell$ is the length of the longest side $L$ of the triangle $T$, and $h$ is the length of the height of the side $L$.
 Remark 1.8.
-We wish to make a comparison between the estimate (1.4) and the estimate (1.5).
-Let $\alpha$ be the angle corresponding to the longest side $L$, and let $m$, $s$, $\ell$ be the lengths of sides $S$, $M$ and $L$ of the triangle $T$, respectively.
-We divide the comparison into two cases:
-Case I. When $\alpha\geq\frac{\pi}{2}$, one has
+Let $\alpha$ be the angle corresponding to the longest side $L$, and here let $s$, $m$, $\ell$ be the lengths of sides $S$, $M$ and $L$ of the triangle $T$, respectively.
+We divide our argument into two cases:
+Case I. When $\alpha\geq\frac{\pi}{2}$ (i.e. $T$ is a right triangle or an obtuse triangle), one has
 $m^{2}+s^{2}\geq 2ms\geq 2ms\sqrt{2}\sin(\alpha+\frac{\pi}{4}),$
@@
 $\ell^{2}\geq 4|T|.$
-This is to say that when $\alpha\geq\frac{\pi}{2}$ (i.e.
-$T$ is a right triangle or an obtuse triangle), the lower bound estimate (1.4) is sharper than (1.5).
-Especially, if $\ell^{2}=4|T|$, then $m=s$ and $\alpha=\pi/4$, which implies that in this setting $T$ should be an isosceles right triangle.
+Especially, if $\ell^{2}=4|T|$, then $m=s$ and $\alpha=\pi/2$, which implies that in this setting $T$ should be an isosceles right triangle.
 Case II.
@@
 $2s^{2}-2s^{2}\cos\alpha<4\times\left(\frac{1}{2}s^{2}\sin\alpha\right),$
-that is, $\ell^{2}<4|T|$. So, in this setting, the lower bound estimate (1.5) is sharper than (1.4).
+that is, $\ell^{2}<4|T|$.
 (2) If furthermore $m\neq s$, then the relation between $\ell^{2}$ and $4|T|$ is uncertain and deeply depends on the choice of values of $m$, $s$.
@@
 $4|T|\approx 23.18,\qquad\ell^{2}\approx 18.7888,$
-which means $4|T|>\ell^{2}$ in this setting; if $m=4$, $s=2$, a simple calculation tells us
-$4|T|\approx 15.454,\qquad\ell^{2}\approx 17.93,$
+which means $4|T|>\ell^{2}$ in this setting. However, if $\ell=4$, $m=3.9823$, $s=2$, a simple calculation tells us
+$4|T|\approx 15.4548,\qquad\ell^{2}=16,$
 which implies $4|T|<\ell^{2}$ in this setting.
-In sum, through the above comparison, readers can see that it is meaningful to give estimates (1.4)-(1.5) for the mixed eigenvalue $\lambda^{MS}_{1}(T)$ simultaneously, and these estimates deeply and intuitively reveal the relation between the eigenvalue $\lambda^{MS}_{1}(T)$ and the length of the longest side, the area of the given triangle $T$.
+In sum, we have:
+  • (a) If $T$ is an obtuse triangle, a right triangle (except the case of isosceles right triangles), or an acute triangle which has the form like the triangle $\ell=4$, $m=3.9823$, $s=2$ constructed in (2) of Case II, then $\ell^{2}>4|T|$.
+  • (b) If $T$ is an isosceles right triangle, then $\ell^{2}=4|T|$.
+  • (c) If $T$ is an isosceles acute triangle or an acute triangle which has the form like the triangle $m=4$, $s=3$, $\alpha=5\pi/12$ constructed in (2) of Case II, then $\ell^{2}<4|T|$.
+On the other hand, it is not hard to see that:
+  • If $0<h<\frac{\ell}{2}$, then $h^{2}<\frac{\ell^{2}}{4}$ and $|T|<\frac{\ell^{2}}{4}$.
+In this situation, $T$ must be a triangle belonging to the category (a) described as above, and moreover, the estimate (1.5) in Theorem 1.7 becomes $\lambda_{1}^{MS}(T)\geq\frac{4\pi^{2}}{\ell^{2}}$.
+Since $|T|=\frac{h\ell}{2}<\frac{\ell^{2}}{4}$, one knows that in this situation, the estimate (1.4) is sharper than the estimate (1.5).
+  • If $h=\frac{\ell}{2}$, then $h^{2}=\frac{\ell^{2}}{4}$ and $|T|=\frac{\ell^{2}}{4}$.
+In this situation, $T$ must be an isosceles right triangle, i.e.
+belonging to the category (b).
+Moreover, the estimate (1.5) becomes $\lambda_{1}^{MS}(T)\geq\frac{\pi^{2}}{|T|}$, which coincides with the lower bound estimate given in (1.4).
+  • If $h>\frac{\ell}{2}$, then $h^{2}>\frac{\ell^{2}}{4}$ and $|T|>\frac{\ell^{2}}{4}$.
+In this situation, $T$ must be a triangle belonging to the category (c), and moreover, the estimate (1.5) becomes $\lambda_{1}^{MS}(T)\geq\frac{\pi^{2}}{h^{2}}$.
+Since $|T|=\frac{h\ell}{2}<h^{2}$, one knows that in this situation, the estimate (1.4) is sharper than the estimate (1.5).
+Therefore, if $T$ is an isosceles right triangle, then the lower bound in the estimate (1.4) or (1.5) coincides with each other, while if $T$ is an obtuse triangle, a right triangle (except the case of isosceles right triangles), or an acute triangle, then the lower bound estimate (1.4) is strictly sharper than the estimate (1.5).
 Given a trapezoid in $\mathbb{R}^{2}$, assume that the mean length of the two parallel sides of this trapezoid is $m$ (we call $m$ the mean width or the average width of the trapezoid), and the height is $h$.
@@
 Let $v$ be the characteristic function corresponding to $\lambda_{1}^{L}(T_{R})$. Construct a function $\phi$ as follows:
-$\phi(x)=\left\{\begin{array}[]{llll}v(x)&\mathrm{if}~x\in T_{R},\\ v(y)&\mathrm{if}~x\in\triangle BCD,\\ v(z)&\mathrm{if}~x\in\triangle ACE,\\ v(w)&\mathrm{if}~x\in\triangle DCE,\end{array}\right$.
+$\phi(x)=\left\{\begin{array}[]{llll}v(x)\qquad&\mathrm{if}~x\in T_{R},\\ v(y)\qquad&\mathrm{if}~x\in\triangle BCD,\\ v(z)\qquad&\mathrm{if}~x\in\triangle ACE,\\ v(w)\qquad&\mathrm{if}~x\in\triangle DCE,\end{array}\right$.
 where $y$ is the symmetric point of $x$ with respect to the side $BC$, $z$ is the symmetric point of $x$ with respect to $AD$, and $w$ is the centrally symmetric point of $x$ with respect to the vertex $C$.
@@
 Let $u$ be a e characteristic function corresponding to the first mixed eigenvalue $\lambda_{1}^{MS}(T)$. Construct a function $\phi$ defined on the rectangle $GBCH$ as follows:
-$\phi=\left\{\begin{array}[]{lll}u&\mathrm{on}~~T,\\ u_{13}&\mathrm{on}~~\triangle GAB,\\ u_{24}&\mathrm{on}~~\triangle AHC,\end{array}\right$.
+$\phi=\left\{\begin{array}[]{lll}u\qquad&\mathrm{on}~~T,\\ u_{13}\qquad&\mathrm{on}~~\triangle GAB,\\ u_{24}\qquad&\mathrm{on}~~\triangle AHC,\end{array}\right$.
 where $u_{13}$ is the function obtained from $u$ by rotating the triangle $\triangle ABD$ with respect to the midpoint of its largest side until it coincides with $\triangle GAB$, and then changing its sign.
@@
 where $\mu_{2}(\mathcal{C})$ is the nonzero Neumann eigenvalue of $\mathcal{C}$.
-By using the basic properties of triangles, one has $0<h\leq\sqrt{3}\ell/2$, and moreover $h=\sqrt{3}\ell/2$ if and only if $T$ is an equilateral triangle.
-This fact, together with (2.7), implies
-$\mu_{2}(\mathcal{C})=\frac{4\pi^{2}}{\ell^{2}}.$
-Due to the definition of $\phi$, it is not hard to see $\int_{\mathcal{C}}\phi=0$, which implies that $\phi$ can be regarded as a trial function for $\mu_{2}(\mathcal{C})$. Hence, we can obtain
+Due to the definition of $\phi$, it is not hard to see $\int_{\mathcal{C}}\phi=0$, which implies that $\phi$ can be regarded as a trial function for $\mu_{2}(\mathcal{C})$.
+Hence, we can obtain
 $\lambda_{1}^{MS}(T)$ $=$ $\frac{\int_{T}|\nabla u|^{2}}{\int_{T}u^{2}}=\frac{\int_{\triangle ABD}|\nabla u|^{2}+\int_{\triangle ACD}|\nabla u|^{2}}{\int_{\triangle ABD}u^{2}+\int_{\triangle ACD}u^{2}}$
@@
 $=$ $\frac{\int_{\mathcal{C}}|\nabla\phi|^{2}}{\int_{\mathcal{C}}\phi^{2}}$
-$\geq$ $\mu_{2}(\mathcal{C})=\frac{4\pi^{2}}{\ell^{2}},$
-which completes the proof of Theorem 1.7. ∎
+$\geq$ $\mu_{2}(\mathcal{C}),$
+which, together with (2.7), completes the proof of Theorem 1.7. ∎
 3 The case of trapezoids
@@
 Therefore, we have
-$\lambda_{k}^{Q^{1}Q^{2}}(\Omega)\int_{\Omega}f^{2}\leq\int_{\Omega}|\nabla f|^{2}=\sum\limits_{j=1}^{k}\alpha_{j}^{2}v_{j}\leq v_{k}\int_{\Omega}f^{2},$
+$\lambda_{k}^{Q^{1}Q^{2}}(\Omega)\int_{\Omega}f^{2}\leq\int_{\Omega}|\nabla f|^{2}=\sum\limits_{j=1}^{k}\alpha_{j}^{2}v_{j}\int_{\Omega}f^{2}\leq v_{k}\int_{\Omega}f^{2},$
 which directly implies
@@
 where $M(x)$ is the inverse function of the function $x=\sum\limits_{j=0}^{\infty}(y-j^{2})_{+}^{\frac{1}{2}}$, $y\geq 0$. Besides, here an abbreviation
-$z_{+}:=\frac{z+|z|}{2}=\left\{\begin{array}[]{ll}z,&when~z\geq 0\\ 0,&when~z<0\end{array}\right$.
+$z_{+}:=\frac{z+|z|}{2}=\left\{\begin{array}[]{ll}z,\quad&when~z\geq 0\\ 0,\quad&when~z<0\end{array}\right$.
 has been used.
@@
 As we know, all the mixed eigenvalues $\lambda_{l}^{Q^{1}Q^{2}}(\Gamma^{\ast})$ of the rectangle $\Gamma^{\ast}$ are
-$\pi^{2}\left(\frac{i^{2}}{m^{2}}+\frac{h^{2}}{b^{2}}\right),\qquad i=1,2,3,\cdots,~~j=0,1,2,\cdots.$
+$\pi^{2}\left(\frac{i^{2}}{m^{2}}+\frac{j^{2}}{h^{2}}\right),\qquad i=1,2,3,\cdots,~~j=0,1,2,\cdots.$
 Together with (3.5), we have
-$\pi^{2}\left(\frac{i^{2}}{m^{2}}+\frac{h^{2}}{b^{2}}\right)=\frac{\pi^{2}}{m^{2}}\left(i^{2}+j^{2}\cdot\frac{m^{2}}{h^{2}}\right)>\frac{\pi^{2}}{m^{2}}\left[i^{2}+j^{2}\cdot(k^{2}-1)\right].$
+$\pi^{2}\left(\frac{i^{2}}{m^{2}}+\frac{j^{2}}{h^{2}}\right)=\frac{\pi^{2}}{m^{2}}\left(i^{2}+j^{2}\cdot\frac{m^{2}}{h^{2}}\right)>\frac{\pi^{2}}{m^{2}}\left[i^{2}+j^{2}\cdot(k^{2}-1)\right].$
 Hence, in our setting here, the $k$-th mixed eigenvalue $\lambda_{k}^{Q^{1}Q^{2}}(\Gamma^{\ast})$ should be attained when $i=k$, $j=0$, that is, $\lambda_{k}^{Q^{1}Q^{2}}(\Gamma^{\ast})=\pi^{2}k^{2}/m^{2}$.
@@
 Let $v$ be the eigenfunction corresponding to the mixed eigenvalue $\lambda_{1}^{l_{1}l_{2}w_{1}}(\Gamma_{R})$, and construct a function $\rho$ as follows
-$\rho(x)=\left\{\begin{array}[]{ll}v(x),&\mathrm{if}~x\in\Gamma_{R},\\ v(y),&\mathrm{if}~x\in I_{s}\setminus\Gamma_{R},\end{array}\right$.
+$\rho(x)=\left\{\begin{array}[]{ll}v(x),\qquad&\mathrm{if}~x\in\Gamma_{R},\\ v(y),\qquad&\mathrm{if}~x\in I_{s}\setminus\Gamma_{R},\end{array}\right$.
 where $y$ is the symmetric point of $x$ with respect to $w_{2}$.
@@
 This research was supported in part by the NSF of China (Grant Nos.
-11801496 and 11926352), the Fok Ying-Tung Education Foundation (China), Hubei Key Laboratory of Applied Mathematics (Hubei University), and Key Laboratory of Intelligent Sensing System and Security (Hubei University), Ministry of Education.
+11801496 and 11926352), the Fok Ying-Tung Education Foundation (China), the Key Project of Jiangxi Provincial Natural Science Foundation (Grant No.
+20252BAC250003), Hubei Key Laboratory of Applied Mathematics (Hubei University), and Key Laboratory of Intelligent Sensing System and Security (Hubei University), Ministry of Education.
+Conflict of interest
+The authors declare that there are no conflicts of interests regarding the publication of this paper.
+Data availability statement
+Data sharing is not applicable to this article as no new data were created or analyzed in this study.